# Leveraging Interpolation Models and Error Bounds for Verifiable Scientific Machine Learning

### v1

@@ 4.1 Validation
Figure 6: Performance with 2-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Predicted error bounds (x-axis) vs. observed errors (y-axis) for each of the 3 interpolation methods
@@
Figure 7: Performance with 4-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Predicted error bounds (x-axis) vs. observed errors (y-axis) for each of the 3 interpolation methods
@@
Figure 8: Performance with 6-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Predicted error bounds (x-axis) vs. observed errors (y-axis) for each of the 3 interpolation methods
@@
Figure 9: Performance with 8-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Predicted error bounds (x-axis) vs. observed errors (y-axis) for each of the 3 interpolation methods
@@
On the other hand, the Delaunay interpolant, TPS RBF, and ReLU MLP all offer surprisingly similar performance on this problem. Additionally, although the error bounds can be an order of magnitude too high in many cases (this is expected since they are bounds), they are generally correlated with the true error. Additionally, the Delaunay bounds can be tight when predicting reasonably low error rates (e.g., relative errors less than 0.5).
@@
For the higher-dimensional embeddings in Figures 7, 8, and 9, the bounds hold but become less tight in higher dimensions for the Delaunay method. In particular, for the $4$-dimensional case, we see that the bounds are asymptotically tight as they approach 00. The bounds remain similar for the TPS RBF in all dimensions, indicating that the embedded training/validation points are tightly clustered in the latent space even for higher-dimensional embeddings, since the TPS RBF bound is purely distance based. On the other hand, the proportion of extrapolation points does increase drastically in high dimensions, which is expected in light of Section 3.4.
@@
  • the $4$-dimensional representation is likely to be optimal since it is the lowest-dimensional latent space where the bounds appear to hold, indicating that the embedding produces a Lipschitz continuous low-noise function over the latent space;
@@
Finally, it is worth noting that for these data sets and train/validate splits, the interpolation methods are an order of magnitude cheaper to train and validate than the ReLU MLP. However, one limitation is that the interpolation methods (especially the Delaunay method) carry a greater proportion amount of their computational cost at inference time, with consequent implications for applications that require real-time predictions. They also require the training data to be held in memory at all times, which can be a privacy concern. Therefore, for many applications they are best suited as a validation technique.
4.2 Testing
For the test set, the distribution of data is clearly different than on the training and validation sets. Only 46% of test points are inside the convex hull of the embedded training data, while 98% of the validation points are inside the convex hull of the embedded training data. Therefore, we separately note the expected and true error rates for test points inside the convex hull vs. the entire test set, to study the difference in performance between in-distribution and out-of-distribution predictions. Results are shown in Table 1. Since we have already performed validation, the results are only shown when using the entire training set (5,553 training points embedded in 4 dimensions).
Method MAE (total) average bound MAE (in hull) average bound (in hull)
@@ 4.2 Testing
Table 1: Averaged error bounds and MAE for all the data (left) and only interpolation points (right) for all 4 methods from Section 2
@@
Compared to errors on the order of $10^{-2}$ obtained when using the full training set during validation in Figure 7, we see much higher relative testing errors across all categories of Table 1 at approximately $0.2$ MAE. This increase in MAE could be attributed to generalization error and increase in out-of-sample predictions. However, upon inspection of the error bounds, we observe that true errors regularly exceed the error bounds in Figure 10. Upon closer inspection, the Delaunay bounds are exceeded by approximately $0.2$ for bounds near 00, which is also the increase in MAE. Adding a fixed amount of $0.2$ to each relative error bound makes the Delaunay bounds asymptotically tight again. This agrees with our hypothesis that the increase is due to generalization error on the test set, which is out-of-sample compared to the training and validation data. We will further explore this in the next section.
@@ 4.3 Interpretability
First notice, that across all three test points, the embeddings are lower-quality compared to in the validation set. This shows that our latent space representation is subject to significant generalization error. However, the results in Figures 11, 12, and 13 paint a nice illustration of the utility of our methods. In particular, Figure 11 (interpolation point that obeys error bounds) shows an embedding that is different but with similar angle-of-attack to the true image, and the training points all closely resemble the embedding. In Figure 12 (interpolation point that violates the error bounds), we see an embedding that has almost no similarity with the true image, but closely matches many training points. In this case, the error in the embedding essentially “tricks” the interpolation bounds. Finally, in Figure 13, the true test image is somewhat similar in shape, but not identical and with different angle-of-attack to its embedding. The embedded image also features an odd distortion on the top right side of the airfoil, which is not present in any of the training images. Thus, it is visually apparent why this point resulted in geometric extrapolation in the latent space.
In all of the above cases, the interpolation methods would not able to account for a poor autoencoder embedding. This likely explains why the generalization error was so great, in particular, exceeding the interpolation error bounds even for the Delaunay method, whereas no predictions in the validation set exceeded the interpolation error bounds for the Delaunay method. In support of the above claim, consider the sample interpolation point and sample extrapolation point from the validation set, shown in Figures 14 and 15, respectively. Note that there is no mention of which points exceed the error bounds here, since on the validation set, every prediction was within the error bounds.
@@
The most striking thing that is immediately obvious from Figures 14 and 15 is that the embedded images are visually much more similar to the true images than in any of Figures 11, 12, or 13. This seems to confirm that it is the failure of our autoencoder embeddings to generalize to the test set, which caused the interpolation error bounds to fail. In this sense, the failure of the interpolation bounds serves is a feature not a bug, as it serves as an alarm bell for an issue in an earlier stage of our scientific machine learning pipeline.
4.4 Summary of Case-Study
In this section, we studied the utility of interpolation methods on a computer-vision-based regression problem from computational physics. We combined common deep learning techniques (representation learning) with interpolation by training a convolutional autoencoder. In this particular example, both the TPS RBF and Delaunay interpolant offered similar performance to the ReLU MLP in both validation and testing. The failure of the GP method indicates that the Gaussian kernel and possibly other assumptions of GP models are not well-suited to this application, most likely due to the smoothness of the embedding.
The advantage to using these interpolation techniques, is that we were able to validate our trained ReLU MLP networks against other methods that do not require any training, giving confidence in their predictions. Furthermore, we made theory-backed choices for the latent space dimension during validation, and quantified the expected errors on the validation domain. On the testing set, we observed significant generalization error. However, we were able to quantify what percentage of this was error was due to the representation $f_{R}(\cdot)$ by leveraging the error bounds of the interpolation methods. Then, by studying the increase in geometric extrapolation in the latent space, we concluded that the majority of testing samples were out-of-distribution for the training samples (at least in the latent space). Finally, leveraging the interpretability of the Delaunay interpolant, we were also able to see that the quality of the latent space embeddings had greatly decreased on the testing set, which is likely a result of this distribution shift.
@@ 5 Discussion and Conclusions
  • other indicators of data set quality such as skew and spacing are less important, and for those methods that are effected, transforming/rescaling the data set (a standard practice) appears to address the issues; and
@@
  • Delaunay interpolation and ReLU MLPs (when optimized) seem to achieve similar performance on real-world problems; and

### v2

@@ 4.1 Validation of latent encoding dimension via interpolation methods
Figure 7: Performance with 2-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Approximation error (x-axis) vs. true error (y-axis) for each of the 3 interpolation methods.
@@
Figure 8: Performance with 4-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Approximation error (x-axis) vs. true error (y-axis) for each of the 3 interpolation methods. See Figure 11 for a zoomed in version of the top right plot.
@@
Figure 9: Performance with 6-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Approximation error (x-axis) vs. true error (y-axis) for each of the 3 interpolation methods.
@@
Figure 10: Performance with 8-dimensional latent space embedding. Top left: MAE and average error bound when predicting lift/drag from geometric shape parameters for various training sizes via interpolation and machine learning methods. Note that results are shown at semilog scale. Bottom left: Only the MAE is shown at linear scale. Right: Approximation error (x-axis) vs. true error (y-axis) for each of the 3 interpolation methods.
@@
On the other hand, the Delaunay interpolant, TPS RBF, and ReLU MLP all offer surprisingly similar performance on this problem. Although the error bounds can be an order of magnitude too high in many cases (this is expected since they are bounds), they are always larger than the true error in MAE and almost always larger at individual points. In particular, we do not observe any points with small approximation error but high true error for these methods. Additionally, the Delaunay bounds can be tight when predicting reasonably low error rates (e.g., relative errors less than 0.5).
@@
For the higher-dimensional embeddings in Figures 8, 9, and 10, the Delaunay bounds hold, with few exceptions. Notably, for the $4$-dimensional case, the bounds appear quite tight near 00, as seen in Figure 11. In higher-dimensional latent spaces, these bounds are less tight.
The bounds for the TPS RBF interpolant remain similar in all dimensions, indicating that the embedded training/validation points are tightly clustered in the latent space even for higher-dimensional embeddings, since the TPS RBF bound is purely distance based. On the other hand, the proportion of extrapolation points does increase drastically in high dimensions, which is expected in light of Section 3.4.
Refer to caption Refer to caption
Figure 11: Left: A zoomed in view of the approximation error (x-axis) vs. true error (y-axis) for the Delaunay interplant on the 4-dimensional latent space embedding (top right plot from Figure 8). The tightness of the bound (and occasional violation) is evident at these small values. Right: Same figure for the 6-dimensional latent space (cf. Figure 9), indicating fewer points with near-zero errors.
@@
  • the $4$-dimensional representation is likely to be optimal since it is the lowest-dimensional latent space where the bounds appear to hold, indicating that the embedding produces a Lipschitz continuous, low-noise function over the latent space;
@@
Finally, it is worth noting that for these data sets and train/validate splits, the interpolation methods are an order of magnitude cheaper to train and validate than the ReLU MLP. However, one limitation is that the interpolation methods (especially the Delaunay method) carry a greater proportion amount of their computational cost at inference time, with consequent implications for applications that require real-time predictions. They also require the training data to be held in memory at all times, which can be a privacy concern. Therefore, for many applications, interpolation methods are best suited as a validation technique.
4.2 Assessing the test set via interpolation methods
We now analyze the 668 points in the “test set” using the autoencoder trained on all the 5553 training data with a 4-dimensional latent space. Looking first at extrapolation prevalence, we find that the geometric distribution of the test data is quite different from the training and validation sets. Only 46% of test points are inside the convex hull of the encoded training data, while 98% of the validation points are inside the convex hull of the encoded training data. To observe whether this geometric difference can account for the performance of interpolation methods deployed on encoded points from the test set, we report MAE and bound results over the “in hull” subset of test points for comparison. Results are shown in Table 1.
All test points In-hull only
Method MAE average bound MAE average bound
@@ 4.2 Assessing the test set via interpolation methods
Table 1: Averaged error bounds and MAE for all the encoded test data (left) and only points inside the convex hull of the training data (right) for all 4 methods from Section 2. These results use the autoencoder trained on all the training data with a 4-dimensional latent space.
@@
Compared to errors on the order of $10^{-2}$ obtained when using the full training set during validation in Figure 8, we see much higher relative testing errors across all categories of Table 1 at approximately $0.2$ MAE. This increase in MAE could be attributed to generalization error and increase in out-of-hull predictions. However, upon inspection of the error bounds, we observe that true errors regularly exceed the error bounds in Figure 12. Upon closer inspection, the Delaunay bounds are exceeded by approximately $0.2$ for bounds near 00, which is also the increase in MAE. Adding a fixed amount of $0.2$ to each relative error bound makes the Delaunay bounds tight again. This agrees with our hypothesis that the increase is due to generalization error on the test set, which is largely encoded outside the convex hull of the training data, compared to the validation data. We will further explore this in the next section.
@@ 4.3 Interpretability enabled by interpolation methods
428324
First notice, that across all three test points, the embeddings are lower-quality compared to those in the validation set (see Figures 16 and 17). This shows that our latent space representation is subject to significant generalization error. However, the results in Figures 13, 14, and 15 paint a nice illustration of the utility of our methods. In particular, Figure 13 (interpolation point that obeys error bounds) shows an embedding that is different but with similar angle-of-attack to the true image, and the training points all closely resemble the embedding. In Figure 14 (interpolation point that violates the error bounds), we see an embedding that has almost no similarity with the true image, but closely matches many training points. In this case, the error in the embedding essentially “tricks” the interpolation bounds. Finally, in Figure 15, the true test image is somewhat similar in shape, but not identical and with different angle-of-attack to its embedding. The embedded image also features an odd distortion on the top right side of the airfoil, which is not present in any of the training images. Thus, it is visually apparent why this point resulted in geometric extrapolation in the latent space.
In all of the above cases, the interpolation methods would not able to account for a poor autoencoder embedding. This likely explains why the generalization error was so great, in particular, exceeding the interpolation error bounds even for the Delaunay method, whereas no predictions in the validation set exceeded the interpolation error bounds for the Delaunay method. In support of the above claim, consider the sample interpolation point and sample extrapolation point from the validation set, shown in Figures 16 and 17, respectively. Note that there is no mention of which points exceed the error bounds here, since on the validation set, every prediction was within the error bounds.
@@
The most striking thing that is immediately obvious from Figures 16 and 17 is that the embedded images are visually much more similar to the true images than in any of Figures 13, 14, or 15. This seems to confirm that it is the failure of our autoencoder embeddings to generalize to the test set, which caused the interpolation error bounds to fail. In this sense, the failure of the interpolation bounds serves is a feature not a bug, as it serves as an alarm bell for an issue in an earlier stage of our scientific machine learning pipeline.
4.4 Summary of case study
In this section, we studied the utility of interpolation methods on a computer-vision-based regression problem from computational physics. We combined common deep learning techniques (representation learning) with interpolation by training a convolutional autoencoder. In this particular example, both the TPS RBF and Delaunay interpolant offered similar performance to the ReLU MLP in both validation and testing. The failure of the GP method indicates that the Gaussian kernel and possibly other assumptions of GP models are not well-suited to this application, most likely due to the smoothness of the embedding. An advantage to using these interpolation techniques is their ability to validate our trained ReLU MLP networks against other methods without any training, giving confidence in the predictions of the neural network.
Furthermore, we made theory-backed choices for the latent space dimension during validation, and quantified the expected errors on the validation set. On the testing set, we observed significant generalization error and computed that more than half the encoded testing data lay outside the convex hull of the encoded training data. We identified that most of the MAE for the true error with the Delaunay interpolant was due to points with near-zero approximation error, and found that adding a fixed amount (0.2) to the Delaunay bound recovered tight bounds and reduced the MAE. Finally, leveraging the interpretability of the Delaunay interpolant, we were also able to see that the quality of the latent space embeddings had greatly decreased on the testing set, which is likely a result of this distribution shift.
Applying our techniques to a problem with latent space dimension $\mathcal{O}\left(10^{1}\right)$ is feasible, provided more data is available, recalling the prevalence of extrapolation for higher dimensions from Figure 5. Interpolation could be more prevalent if the data is chosen in a suitably intentional fashion (instead of randomly) or if the query domain of interest is largely contained in the convex hull of the data. If the requisite latent space dimension is $\mathcal{O}\left(10^{2}\right)$ or larger, interpolation methods are unlikely to provide significant insight.
@@ 5 Discussion and Conclusions
  • other indicators of data set quality such as skew and spacing are less important, and for those methods that are affected, transforming/rescaling the data set (a standard practice) appears to address the issues; and
@@
  • Delaunay interpolation and ReLU MLPs (when optimized) seem to achieve similar performance on real-world problems;